# Camera-based 3D Semantic Scene Completion with Sparse Guidance Network

### v1

@@ Abstract
Semantic scene completion (SSC) aims to predict the semantic occupancy of each voxel in the entire 3D scene from limited observations, which is an emerging and critical task for autonomous driving. Recently, many studies have turned to camera-based SSC solutions due to the richer visual cues and cost-effectiveness of cameras. However, existing methods usually rely on sophisticated and heavy 3D models to directly process the lifted 3D features that are not discriminative enough for clear segmentation boundaries. In this paper, we adopt the dense-sparse-dense design and propose an end-to-end camera-based SSC framework, termed SGN, to diffuse semantics from the semantic- and occupancy-aware seed voxels to the whole scene based on geometry prior and occupancy information. By designing hybrid guidance (sparse semantic and geometry guidance) and effective voxel aggregation for spatial occupancy and geometry priors, we enhance the feature separation between different categories and expedite the convergence of semantic diffusion. Extensive experimental results on the SemanticKITTI dataset demonstrate the superiority of our SGN over existing state-of-the-art methods. Code is available at https://github.com/Jieqianyu/SGN.
Introduction
@@
3D scene understanding constitutes the bedrock of autonomous driving systems. By precisely perceiving the occupancy and semantics of their surroundings, autonomous vehicles can make informed decisions and navigate safely. To this end, Semantic Scene Completion (SSC) has been introduced to predict the semantic occupancy of each voxel of the entire 3D scene from limited observation. However, accurately estimating the semantics and geometry of the real world from partial observations is challenging due to the complexities presented by real-world scenarios.
@@
Figure 1: (a) Fully dense processing with heavy and complex 3D model. (b) MAE-like architecture in a “sparse-to-dense” manner. (c) Our “dense-sparse-dense” design with hybrid guidance and semantic diffusion.
@@
SSC has attracted extensive studies due to its application prospects for downstream tasks such as mapping and planning. When working with outdoor driving scenes, LiDAR has emerged as a popular input modality for many existing methods Roldão, de Charette, and Verroust-Blondet (2020); Cheng et al. (2021); Rist et al. (2021); Yan et al. (2021); Xia et al. (2023); Li et al. (2023b) to capture 3D information of surroundings, but it suffers from high-cost sensors. Recently, there has been a shift towards camera-based SSC solutions. As the pioneer, MonoScene Cao and de Charette (2022) proposed the first framework for monocular 3D SSC, utilizing mapping projection to lift RGB images to 3D volumes processed with the 3D UNet. Afterward, many camera-based methods such as OccDepth Miao et al. (2023), SurroundOcc Wei et al. (2023), and OccFormer Zhang, Zhu, and Du (2023) are developed with a similar pipeline consisting of the image backbone, view transformer, and 3D model, as illustrated in Fig.1 (a). However, they rely on sophisticated and heavy 3D models to directly process the lifted 3D features that are not discriminative enough for clear segmentation boundaries. We explain that the lifted 3D features by 3D-2D mapping projection Cao and de Charette (2022) contain many ambiguities due to the assumption of the uniform depth distribution and 2D-3D methods such as LSS Philion and Fidler (2020) only utilize coarse surface information from depth distribution estimation.
@@
On the other hand, VoxFormer Li et al. (2023d) proposed an MAE-like architecture to complete non-visible structures using constructed visible areas. It adopts the two-stage framework, with the first stage for query proposal and the second stage for densification and segmentation. By completing the 3D scene in a sparse-to-dense manner shown in Fig.1 (b), VoxFormer is more efficient and scalable than the dense processing with heavy 3D models mentioned above. However, it still suffers from several limitations. Firstly, the densification stage is mainly considered from the perspective of scene completion based on queries. The intra-category feature separation of queries is neglected. Besides, the second stage only considers the information from the queries that only include partial observation and are not always accurate, increasing the difficulty of subsequent completion and segmentation. Finally, the two-stage training and inference cannot fully consider global information due to the independent optimization of different stages. The spatial occupancy information from the first stage is also not fully utilized.
@@
To address the above problems, we propose a novel end-to-end camera-based SSC framework, Sparse Guidance Network (SGN), to diffuse semantics from the semantic- and occupancy-aware seed voxels to the whole scene based on geometry prior and occupancy information, as illustrated in Fig.1 (c). Specifically, we employ the dense-sparse-dense design to implement the semantic diffusion of seed features, avoiding relying on heavy 3D models to process coarse scene representations that are not discriminative enough. Firstly, to dynamically select sparse seed voxels and provide occupancy-aware information, we redesign the sparse voxel proposal network to directly process points generated by depth prediction with the coarse-to-fine paradigm. And by further designing hybrid guidance (sparse semantic and geometry guidance) and effective voxel aggregation for spatial occupancy and geometry priors, we enhance the intra-category feature separation and expedite the convergence of the semantic diffusion. We also devise the multi-scale semantic diffusion module using anisotropic convolutions Li et al. (2020) for flexible receptive filed while reducing the computation resources. By this means, our end-to-end SGN is lightweight while having a more powerful representation ability.
@@
Extensive experiments on the challenging SemanticKITTI dataset Behley et al. (2019) demonstrate the superiority of our SGN over existing state-of-the-art methods. For example, on the validation set, even our light version SGN-L achieves notable scores of 14.80% mIoU and 45.45% IoU with only 12.5 M parameters and 7.16 G memory for training, exceeding VoxFormer-T by 1.45% points in mIoU and 1.30% points in IoU while being more lightweight and less memory consumption.
@@
  • We propose an end-to-end camera-based SSC framework called SGN, propagating semantics from the semantic- and occupancy-aware seed voxels to the whole scene based on geometry prior and occupancy information.
@@
  • We adopt the dense-sparse-dense design and propose hybrid guidance and effective voxel aggregation to enhance intra-categories feature separation and expedite the convergence of the semantic diffusion.
@@
  • Extensive experiments on the SemanticKITTI benchmark demonstrate the effectiveness of our SGN, which is more lightweight and achieves the new state-of-the-art.
Related Works
Semantic Scene Completion.
@@
Due to the vital application of semantic occupancy prediction in autonomous driving, SSC has attracted extensive attention. After the release of the large-scale outdoor benchmark SemanticKITTI Behley et al. (2019), many outdoor SSC methods have emerged. According to the input modality, existing outdoor methods can be mainly classified into LiDAR-based and camera-based methods. LiDAR-based methods consider LiDAR a primary modality to enable accurate 3D semantic occupancy prediction. Following the pioneering SSCNet Song et al. (2017), UDNet Zou et al. (2021) exploits a single 3D U-Net framework to obtain predictions from the grids generated by the LiDAR points, resulting in extra computation overhead of empty voxels. Afterward, LMSCNet Roldão, de Charette, and Verroust-Blondet (2020) introduces the 2D CNN for feature encoding, and SGCNet Zhang et al. (2018) uses spatial group convolutions to improve efficiency. Some solutions focus on multi-view fusion Cheng et al. (2021), local implicit functions Rist et al. (2021), and knowledge distillation Xia et al. (2023) for SSC. Besides, the relationships between semantic segmentation and scene completion are explored. For example, JS3C-Net Yan et al. (2021) and SSA-SC Yang et al. (2021) design a semantic segmentation network to assist the semantic scene completion. SSC-RS Mei et al. (2023) design multi-branch network to fuse semantic and geometry features hierarchically.
@@
Figure 2: Overall framework of our SGN. The image encoder extracts 2D features to provide the foundation for 3D features lifted by the view transformation. Then auxiliary occupancy head is applied to provide geometry guidance. Before sparse semantic guidance, depth-based occupancy prediction is utilized for voxel proposals of indexing seed features. Afterward, the voxel aggregation layer forms the informative voxel features processed by the multi-scale semantic diffusion for the final semantic occupancy prediction. $\mathrm{KT}$ denotes the knowledge transfer layer for geometry prior.
@@
Recently, camera-based perception such as detection Huang et al. (2021); Wang et al. (2022); Liu et al. (2022); Li et al. (2022); Wang et al. (2021); Chen et al. (2022b) and segmentation Li et al. (2022); Zhou and Krähenbühl (2022); Chen et al. (2022a) is currently more attractive due to cameras’ richer visual cues and cost-effectiveness. And there is also a shift towards camera-based solutions Cao and de Charette (2022); Huang et al. (2023) to SSC. MonoScene Cao and de Charette (2022) first proposed to infer 3D SSC from a single monocular RGB image, which applied a classical 3D UNet network to process the voxel features projected along the line of sight. Afterward, TPVFormer Huang et al. (2023) proposed a tri-perspective view (TPV) representation to describe the fine-grained 3D structure of a scene. VoxFormer Li et al. (2023d) proposed an MAE-like architecture to complete non-visible structures using constructed visible areas. OccFormer Zhang, Zhu, and Du (2023) designed a dual-path transformer network. And SurroundOcc Wei et al. (2023) applied 3D convolutions to upsample multi-scale voxel features progressively and devised a pipeline to generate dense SSC ground truth. Symphonize Jiang et al. (2023) modeled the scene volume with a sparse set of instance queries with context awareness. Some methods Miao et al. (2023); Li et al. (2023a) leveraged implicit stereo depth information and stereo matching to resolve geometry ambiguity. Besides, there are some SSC solutions Huang et al. (2023); Wei et al. (2023); Gan et al. (2023); Li et al. (2023e) for multi-view cameras. And multiple benchmarks Li et al. (2023c); Tian et al. (2023); Wang et al. (2023) are proposed to facilitate the SSC’s development.
@@
We focus on camera-based SSC in outdoor scenarios. Compared with the existing works, our SGN proposes to diffuse semantics from the semantic- and occupancy-aware seed voxels to the whole scene based on the geometry prior and occupancy information. SGN avoids relying on heavy and sophisticated 3D models to handle lifted voxel features with rough geometry context like many existing SSC methods Cao and de Charette (2022); Wei et al. (2023); Zhang, Zhu, and Du (2023); Miao et al. (2023); Jiang et al. (2023). Our method is built on the recent two-stage method VoxFormer Li et al. (2023d). However, unlike VoxFormer, our SGN is end-to-end, which adopts the dense-sparse-dense design and proposes hybrid guidance and effective voxel aggregation to enhance intra-categories feature separation and expedite the convergence of the semantic diffusion. Compared with VoxFormer, our SGN achieved better performance while being more lightweight and requiring less memory consumption.
Method
Overview
@@
We show the overall framework of our SGN in Fig.2. SGN adopts the dense-sparse-dense design and propagates semantics from the semantic- and occupancy-aware seed voxels to the whole scene based on geometry prior and occupancy information. SGN takes RGB images as the input and extracts 2D features using the image encoder. Then the 3D features are obtained through the view transformation. For dynamically indexing the seed voxels, we generate the sparse voxel proposals based on depth prediction. Then according to the voxel proposals and 3D features, the hybrid guidance is designed to inject semantic and geometry cues and facilitate feature learning. Furthermore, we develop the voxel aggregation layer to form the informative voxel features, which are further processed by our multi-scale semantic diffusion module for the final semantic occupancy prediction.
Image Encoder.
@@
We use ResNet-50 He et al. (2016) with FPN Lin et al. (2017) to construct the image encoder for extracting 2D features from RGB images. The extracted features $\textbf{F}^{2D}\in\mathbb{R}^{N_{t}\times C\times H\times W}$ provide a strong foundation for the subsequent voxel features, where $N_{t}$ is the image number of temporal inputs, $C$ is the feature channel and $(H,W)$ denotes the image resolution.
View Transformation.
@@
Similar to MonoScene Cao and de Charette (2022), we construct 3D features by sampling 2D features via 3D-2D projection mapping with camera parameters. The simple projection mapping operation provides coarse volumetric scene representation for the latter contextual modeling. And it is more convenient and concise than learnable LSS Philion and Fidler (2020) and cross-attention mechanism Li et al. (2022).
@@
Let $\textbf{x}\in\mathbb{R}^{X\times Y\times Z\times 3}$ denote the centroid of $X\times Y\times Z$ voxels in world coordinates. We establish the projection mapping $\pi(\textbf{x})$ using the camera parameters $(K,T)$, where $K$ and $T$ are the cameras intrinsic and extrinsic matrices directly provided in KITTI Geiger, Lenz, and Urtasun (2012). We take the average of sampled features from different images for each voxel. And the features of voxels outside the field of view (FOV) are set to zero. Mathematically, the 3D features $\textbf{F}^{3D}\in\mathbb{R}^{C\times X\times Y\times Z}$ are sampled from the 2D features $\textbf{F}^{2D}$ as follows:
$\textbf{F}^{3D}=W\cdot\sum_{t=1}^{N_{t}}{[\phi_{\pi(\textbf{x})}(\textbf{F}_{t}^{2D})\cdot M_{t}^{FOV}]}$ (1)
@@
where $\phi_{a}(b)$ is the sampling function that samples features $b$ at coordinates $a$, $\textbf{F}^{2D}_{t}$ is the 2D features of image $\textbf{I}_{t}$, $M^{FOV}_{t}\in\mathbb{R}^{1\times H\times W}$ is the binary mask indicating the field of view of image $\textbf{I}_{t}$, $\delta_{p}$ is the number of hit images for point $p$ in x, $W_{p}$ is the weight value for points $p$ in $W$.
Feature Learning with Hybrid Guidance
As discussed above, most existing methods design heavy models to directly process the 3D features $\textbf{F}^{3D}$ produced by the view transformation module for the final semantic scene prediction. We argue that the coarse scene representation $\textbf{F}^{3D}$ is not discriminative enough to get clear segmentation boundaries, which slows down the convergence of the model. Therefore, we propose sparse semantic guidance and geometry guidance to inject semantic and geometry cues for informative voxel features.
Geometry Guidance.
We first attach the auxiliary 3D occupancy head as the geometry guidance after the 3D features from the view transformation module to provide coarse geometry awareness. Specifically, we construct the 3D occupancy head with an anisotropic convolution layer Li et al. (2020) and a linear layer. In the spirit of Zhou et al. (2020), the anisotropic convolution decomposes a 3D convolution operation into three consecutive 1D convolutions in different directions. Additionally, each of these 1D convolutions is equipped with a mixer containing distinct kernel sizes, enhancing the model’s ability to learn and extract meaningful features from the input data. It can provide flexible receptive fields while alleviating resource consumption. By predicting the 3D occupancy $\hat{\textbf{Y}}_{o}$ on the lifted 3D features $\textbf{F}^{3D}$ using the auxiliary head, we apply the guidance on the coarse scene representation and provide the geometry prior for the latter seed features’ semantic prediction and diffusion. We optimize the occupancy probability with binary cross-entropy loss:

### v2

@@ Abstract
Semantic scene completion (SSC) aims to predict the semantic occupancy of each voxel in the entire 3D scene from limited observations, which is an emerging and critical task for autonomous driving. Recently, many studies have turned to camera-based SSC solutions due to the richer visual cues and cost-effectiveness of cameras. However, existing methods usually rely on sophisticated and heavy 3D models to process the lifted 3D features directly, which are not discriminative enough for clear segmentation boundaries. In this paper, we adopt the dense-sparse-dense design and propose a one-stage camera-based SSC framework, termed SGN, to propagate semantics from the semantic-aware seed voxels to the whole scene based on spatial geometry cues. Firstly, to exploit depth-aware context and dynamically select sparse seed voxels, we redesign the sparse voxel proposal network to process points generated by depth prediction directly with the coarse-to-fine paradigm. Furthermore, by designing hybrid guidance (sparse semantic and geometry guidance) and effective voxel aggregation for spatial geometry cues, we enhance the feature separation between different categories and expedite the convergence of semantic propagation. Finally, we devise the multi-scale semantic propagation module for flexible receptive fields while reducing the computation resources. Extensive experimental results on the SemanticKITTI and SSCBench-KITTI-360 datasets demonstrate the superiority of our SGN over existing state-of-the-art methods. And even our lightweight version SGN-L achieves notable scores of 14.80% mIoU and 45.45% IoU on SeamnticKITTI validation with only 12.5 M parameters and 7.16 G training memory. Code is available at https://github.com/Jieqianyu/SGN.
Index Terms:
Semantic Scene Completion, Sparse Guidance Network, Hybrid Guidance, Voxel Aggregation.
I Introduction
@@
In recent years, there has been significant attention and rapid progress in 3D scene understanding, which constitutes the bedrock of autonomous driving systems and robotics. By precisely perceiving the occupancy and semantics of their surroundings, autonomous vehicles, and robotics can make informed decisions and navigate safely. To this end, Semantic Scene Completion (SSC) has been introduced to predict the semantic occupancy of each voxel of the entire 3D scene from limited observation. SSC helps create a more comprehensive representation of the environment, which includes filling in the gaps or missing information in the sensor data. This can be essential for agents to identify obstacles, understand the road layout, and make safe decisions. However, accurately estimating the semantics and geometry of the real world from partial observations is challenging due to the complexities presented by real-world scenarios.
@@
Figure 1: (a) Fully dense processing with heavy and complex 3D model. (b) MAE-like architecture in a “sparse-to-dense” manner. (c) Our “dense-sparse-dense” design with hybrid guidance and semantic propagation. “A” denotes the voxel aggregation layer for geometry cues.
@@
SSC has attracted extensive studies due to its application prospects for downstream tasks such as mapping and planning. When working with outdoor driving scenes, LiDAR has emerged as a popular input modality for many existing methods [1, 2, 3, 4, 5, 6] to capture 3D information of surroundings, but it suffers from high-cost sensors. Recently, there has been a shift towards camera-based SSC solutions. As the pioneer, MonoScene [7] proposed the first framework for monocular 3D SSC, utilizing mapping projection to lift RGB images to 3D volumes processed with the 3D UNet. Afterward, many camera-based methods such as OccDepth [8], SurroundOcc [9], and OccFormer [10] are developed with a similar pipeline consisting of the image backbone, view transformer, and 3D model, as illustrated in Figure 1 (a). However, they rely on sophisticated and heavy 3D models to process the lifted 3D features directly, which are not discriminative enough for clear segmentation boundaries. We explain that the lifted 3D features by 3D-2D mapping projection [7] contain many ambiguities due to the assumption of the uniform depth distribution and 2D-3D methods such as LSS [11] only utilize coarse surface information from depth distribution estimation.
@@
On the other hand, VoxFormer [12] proposed an MAE-like architecture to complete non-visible structures using constructed visible areas. It adopts the two-stage framework, with the first stage for query proposal and the second stage for densification and segmentation. By completing the 3D scene in a sparse-to-dense manner shown in Figure 1 (b), VoxFormer is more efficient and scalable than the dense processing with complicated 3D models mentioned above. However, it still suffers from several limitations. Firstly, the densification stage is mainly considered from the perspective of scene completion based on queries. The intra-category feature separation of queries is neglected. Besides, the second stage only considers the information from the queries that only include partial observation and are not always accurate, increasing the difficulty of subsequent completion and segmentation. Finally, the two-stage training and inference cannot fully consider global information due to the independent optimization of different stages. The geometry information from the first stage is also not fully utilized.
@@
To address the above problems, we propose a novel one-stage camera-based SSC framework, Sparse Guidance Network (SGN), to propagate semantics from the semantic-aware seed voxels to the whole scene based on spatial geometry cues, as illustrated in Figure 1 (c). Specifically, we employ the dense-sparse-dense design to implement the semantic propagation of semantic-aware seed features, avoiding relying on heavy 3D models to process coarse scene representations that are not discriminative enough. Firstly, to dynamically select sparse seed voxels and encode depth-aware context, we redesign the sparse voxel proposal network to directly process points generated by depth prediction with the coarse-to-fine paradigm. And by further designing hybrid guidance (sparse semantic and geometry guidance) and effective voxel aggregation for spatial geometry cues, we enhance the intra-category feature separation and expedite the convergence of the semantic propagation. We also devise the multi-scale semantic propagation module using anisotropic convolutions [13] for flexible receptive fields while reducing the computation resources. By this means, our SGN is lightweight while having a more powerful representation ability.
@@
Extensive experiments on the challenging SemanticKITTI [14] and SSCBench-KITTI-360 [15] datasets demonstrate the superiority of our SGN over existing state-of-the-art methods. For example, on the SemanticKITTI validation set, even our lightweight version SGN-L achieves notable scores of 14.80% mIoU and 45.45% IoU with only 12.5 M parameters and 7.16 G memory for training, exceeding VoxFormer-T by 1.45% points in mIoU and 1.30% points in IoU while being more lightweight and less memory consumption.
@@
  • We propose a one-stage camera-based SSC framework called SGN, propagating semantics from the semantic- and occupancy-aware seed voxels to the whole scene based on spatial geometry cues.
@@
  • We adopt the dense-sparse-dense design and propose hybrid guidance and effective voxel aggregation to enhance intra-categories feature separation and expedite the convergence of the semantic propagation.
@@
  • Extensive experiments on the SemanticKITTI and SSCBench-KITTI-360 benchmarks demonstrate the effectiveness of our SGN, which is more lightweight and achieves the new state-of-the-art.
II Related Works
Semantic Scene Completion
Due to the vital application of semantic occupancy prediction in autonomous driving, SSC has attracted extensive attention. After the release of the large-scale outdoor benchmark SemanticKITTI [14], many outdoor SSC methods have emerged. According to the input modality, existing outdoor methods can be mainly classified into LiDAR-based and camera-based methods.
@@ II Related Works
LiDAR-based methods consider LiDAR a primary modality to enable accurate 3D semantic occupancy prediction. Following the pioneering SSCNet [16], UDNet [17] exploits a single 3D U-Net framework to obtain predictions from the grids generated by the LiDAR points, resulting in extra computation overhead of empty voxels. Afterward, LMSCNet [1] introduces the 2D CNN for feature encoding, and SGCNet [18] uses spatial group convolutions to improve efficiency. Some solutions focus on multi-view fusion [2], local implicit functions [3], and knowledge distillation [5] for SSC. Besides, the relationships between semantic segmentation and scene completion are explored. For example, JS3C-Net [4] and SSA-SC [19] design a semantic segmentation network to assist the semantic scene completion. SSC-RS [20] design multi-branch network to fuse semantic and geometry features hierarchically.
@@
Figure 2: Overall framework of our SGN. The image encoder extracts 2D features, establishing the foundation for the 3D features generated through view transformation. An auxiliary occupancy head is applied to provide geometry guidance. The sparse semantic guidance consists of two parts: sparse voxel proposal and semantic guidance. The depth-based occupancy prediction is designed for the sparse voxel proposal. This proposal, along with the 3D features, is fed into the subsequent semantic guidance (depicted in Figure 3) to index seed features and inject semantic context into these seed features. Afterward, the voxel aggregation layer combines the semantic-aware seed features, geometry prior from the non-seed features, and occupancy-aware features from the depth-based occupancy prediction. This forms the informative voxel features processed by the multi-scale semantic propagation for the final prediction.
@@
Camera-based methods. Recently, camera-based perception such as detection [21, 22, 23, 24, 25, 26] and segmentation [24, 27, 28] is currently more attractive due to cameras’ richer visual cues and cost-effectiveness. And there is also a shift towards camera-based solutions [7, 29] to SSC. MonoScene [7] first proposed to infer 3D SSC from a single monocular RGB image, which applied a classical 3D UNet network to process the voxel features projected along the line of sight. Afterward, TPVFormer [29] proposed a tri-perspective view (TPV) representation to describe the fine-grained 3D structure of a scene. VoxFormer [12] proposed an MAE-like architecture to complete non-visible structures using constructed visible areas. OccFormer [10] designed a dual-path transformer network. And SurroundOcc [9] applied 3D convolutions to upsample multi-scale voxel features progressively and devised a pipeline to generate dense SSC ground truth. Symphonize [30] modeled the scene volume with a sparse set of instance queries with context awareness. Some methods [8, 31] leveraged implicit stereo depth information and stereo matching to resolve geometry ambiguity. NDC-scene [32] extends the 2D feature map to a Normalized Device Coordinates (NDC) space to alleviate the feature ambiguity, pose ambiguity and computation imbalance. Besides, there are some SSC solutions [29, 9, 33, 34] for multi-view cameras. And multiple benchmarks [15, 35, 36] are proposed to facilitate the SSC’s development.
@@
We focus on camera-based SSC in outdoor scenarios. Compared with the existing works, our SGN proposes to propagate semantics from the semantic-aware seed voxels to the whole scene based on spatial geometry cues. SGN avoids relying on heavy and sophisticated 3D models to handle lifted voxel features with rough geometry context like many existing SSC methods [7, 9, 10, 8, 30]. Our method is built on the recent two-stage method VoxFormer [12]. However, unlike VoxFormer, our SGN is one-stage, which adopts the dense-sparse-dense design and proposes hybrid guidance and effective voxel aggregation to enhance intra-categories feature separation and expedite the convergence of the semantic propagation. Compared with VoxFormer, our SGN achieved better performance while being more lightweight and requiring less memory consumption.
Camera-based 3D Perception
Camera-based 3D perception, encompassing domains such as 3D detection [21, 22, 23, 24, 25, 37, 38, 26] and segmentation [24, 27, 28, 39, 40], has gained increasing traction owing to the rich visual cues provided by cameras and their cost-effectiveness. Various monocular-based approaches have adapted 2D techniques to the 3D domain, such as FCOS3D [25] and DETR3D [22]. In recent times, a significant shift has been observed in camera-based research toward Bird’s Eye View (BEV) representations [21, 24, 41, 42, 43, 27, 44, 39], facilitated by view transformation techniques such as LSS [11], OFT [45], and the cross-attention module [24]. For example, BEVDet [21] and BEVDepth [41] incorporate depth estimation to facilitate the transformation from perspective view to BEV. Additionally, BEVFormer [24] employs cross-attention to inject cues from image features to BEV queries effectively. The efficacy of BEV-based perception [24, 41, 43, 46, 20, 47] has been validated by these advancements. However, for Semantic Scene Completion (SSC) tasks, the utilization of 3D voxel representations, which encapsulate more volumetric information, becomes imperative. As such, the quest to devise discriminative 3D scene representations and to process voxel features both effectively and efficiently remains a vibrant area of ongoing research and exploration.
III Method
III-A Overview
@@
We show the overall framework of our SGN in Figure 2. SGN adopts the dense-sparse-dense design and propagates semantics from the semantic-aware seed voxels to the whole scene based on spatial geometry cues from the non-seed features and features from the depth-based occupancy prediction. SGN takes RGB images as the input and extracts 2D features using the image encoder. Then the 3D features are obtained through the view transformation. For dynamically indexing seed voxels, we generate the sparse voxel proposal based on depth prediction. Then according to the proposal and 3D features, the hybrid guidance is designed to inject semantic and geometry cues and facilitate feature learning. Furthermore, we develop the voxel aggregation layer to form the informative voxel features, which are processed by our multi-scale semantic propagation module for the final semantic occupancy prediction.
Image Encoder
@@
We use ResNet-50 [48] with FPN [49] to construct the image encoder for extracting 2D features from RGB images. The extracted features $\textbf{F}^{2D}\in\mathbb{R}^{N_{t}\times C\times H\times W}$ provide a strong foundation for the subsequent voxel features, where $N_{t}$ is the image number of temporal inputs, $C$ is the feature channel and $(H,W)$ denotes the image resolution.
View Transformation
@@
Similar to MonoScene [7], we construct 3D features by sampling 2D features via 3D-2D projection mapping with camera parameters. The simple projection mapping operation provides coarse volumetric scene representation for the latter contextual modeling. And it is more convenient and concise than learnable LSS [11] and cross-attention mechanism [24].
[…]
$[x_{c},y_{c},z_{c}]^{T}=\textbf{R}\cdot p+\textbf{t}$ (1)
$z_{c}\circ[u,v,1]^{T}=\textbf{K}\cdot[x_{c},y_{c},z_{c}]^{T}$ (2)
where $\circ$ denote element-wise product.
@@
We take the average of sampled features from different images for each voxel. And the features of voxels outside the field of view (FOV) are set to zero. Mathematically, the 3D features $\textbf{F}^{3D}\in\mathbb{R}^{C\times X\times Y\times Z}$ are sampled from the 2D features $\textbf{F}^{2D}$ as follows:
$\textbf{F}^{3D}=W\cdot\sum_{t=1}^{N_{t}}{[\phi_{\pi(\textbf{x})}(\textbf{F}_{t}^{2D})\cdot\textbf{M}_{t}^{FOV}]}$ (3)
@@
where $\phi_{a}(b)$ is the sampling function that samples features $b$ at coordinates $a$, $\textbf{F}^{2D}_{t}$ is the 2D features of image $\textbf{I}_{t}$, $\textbf{M}^{FOV}_{t}\in\mathbb{R}^{1\times X\times Y\times Z}$ is the binary mask indicating the field of view of image $\textbf{I}_{t}$, $\delta_{p}$ is the number of hit images for point $p$ in x, $W_{p}$ is the weight value for points $p$ in $W$.
III-B Feature Learning with Hybrid Guidance
As discussed above, most existing methods design heavy and complicated models to directly process the 3D features $\textbf{F}^{3D}$ produced by the view transformation module for the final semantic scene prediction. We argue that the coarse scene representation $\textbf{F}^{3D}$ is not discriminative enough to get clear segmentation boundaries, which slows down the convergence of the model. Therefore, we propose sparse semantic guidance and geometry guidance to inject semantic and geometry cues for informative voxel features.
III-B1 Geometry Guidance.
We first attach the auxiliary 3D occupancy head as the geometry guidance after the 3D features from the view transformation module to provide coarse geometry awareness. Specifically, we construct the 3D occupancy head with an anisotropic convolution layer [13] and a linear layer. In the spirit of [51], the anisotropic convolution decomposes a 3D convolution operation into three consecutive 1D convolutions in different directions. Additionally, each of these 1D convolutions is equipped with a mixer containing distinct kernel sizes, enhancing the model’s ability to learn and extract meaningful features from the input data. It can provide flexible receptive fields while alleviating resource consumption. By predicting the 3D occupancy $\hat{\textbf{Y}}_{o}$ on the lifted 3D features $\textbf{F}^{3D}$ using the auxiliary head, we apply the guidance on the coarse scene representation and provide the geometry prior for the latter seed features’ semantic prediction and propagation. We optimize the occupancy probability with binary cross-entropy loss: